ROTATIONAL SPECTRA OF C_{2}D_{4} − H_{2}S, C_{2}D_{4} − D_{2}S, C_{2}D_{4} − HDS and ^{13}CCH_{4} − H_{2}S COMPLEXES

MAUSUMI GOSWAMI and \underline{E. ARUNAN}, Department of Inorganic and Physical Chemistry, Indian Institute of Science, Bangalore 560012, India.

The rotational spectra of C_{2}D_{4} − H_{2}S, C_{2}D_{4} − HDS, C_{2}D_{4} − D_{2}S and ^{13}CCH_{4} − H_{2}S isotopomers have been measured. The spectra for C_{2}D_{4} − H_{2}S, C_{2}D_{4} − D_{2}S and ^{13}CCH_{4} − H_{2}S show four line pattern as was observed for the parent species [1]. C_{2}D_{4} − HDS shows doubling of transitions. These data combined with our earlier data of C_{2}H_{4} − H_{2}S, C_{2}H_{4} − HDS, C_{2}H_{4} − H_{2}^{34}S and C_{2}H_{4} − D_{2}S confirm that the smaller splitting (0.14 MHz in (B + C)/2 for the parent isotopomer) arises from the tunneling of ethylene hydrogens whereas the larger spilitting (1.67 MHz in (B + C)/2 for the parent isotopomer) arises from the interchange of equivalent hydrogens of H_{2}S. The splittings observed for various isotopic substitutions confirm that the dynamics of both these motions involve both ethylene and hydrogen-sulphide. Molecular symmetry group analysis has been performed where both ethylene and hydrogen motions have been considered. The experimental intensity patterns are consistent with the calculated statistical weights.

1. E. Arunan and co-workers, Chem. Phys. Lett., 393, 2004, 22.